Circular No. 8038

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002ka B. F. Madore, Carnegie Observatories, reports the discovery of a su-

pernova on a spectrogram (range 400–800 nm) taken Dec. 28.256 UT with the Baade 6.5-m telescope, showing the object to be of type Ia, ≈ 26 days past maximum light. The telescope-pointing position was α = 9^{h}30^{m}42^{s}.0, δ = −14^{o}47′04′′ (equinox 2000.0). Following requests by the Central Bureau, additional information has been received from R. Kushida (Yatsugatake South Base Observatory; via S. Nakano, Sumoto, Japan) and from W. Li (University of California, Berkeley). An image taken by Kushida (measured by Y. Kushida) on Dec. 28.708 yields mag 17.5 and the following precise position: α = 9^{h}30^{m}41^{s}.62, δ = −14^{o}46′55′′.3, which is 2′′.5 west and and 1′′.7 north of the center of a small galaxy. SN 2002ka is also present on KAIT images (cf. IAUC 7906) at position end figures 41^{s}.64, 55′′.2, or 2′′.4 west and 2′′.0 north of the diffuse nucleus of the host galaxy; Li also provides the following magnitudes for the supernova: Apr. 8.3, [19:; Nov. 19.5, 14.8; 30.5, 15.0; Dec. 11.5, 15.3.

SUPERNOVAE 2002kb, 2002kc, 2002kd, 2002ke Further to IAUC 7981, L. Strolger, Space Telescope Science Institute,

on behalf of the GOODS Treasury Team (+ P. Garnavich, C. Kretchmer, and E. Taylor) and the Hubble Higher-z Supernova Team, reports the discovery of four new apparent supernovae on images of the Chandra Deep Field South, each detected by subtracting from ACS images of Dec. 19– 22 those taken on July 31–Aug. 4, Sept. 19–22, and Oct. 31–Nov. 3. SN 2002kb was not detected in the Dec. 19–22 images, but was clearly visible earlier; the remaining objects were detected independently in each of four separate z-band (F850lp) exposures, as well as in the two i-band (F775W) and the two v-band (F606W) exposures. The offsets given below are from the centers of the nearest visible galaxies. Further z-band magnitudes: SN 2002kb, Sept. 22.16 UT, 24.1; Oct. 30.82, 24.2; Dec. 20.10, [26.5. 2002kc, Nov. 3.41, [26. SN 2002kd, Nov. 1.00, [28. SN 2002ke, Sept. 20.30, [27.5.

SN  2002 UT  α_{2000}  δ_{2000}  z  Offset
2002kb July 31.74 3^{h}32^{m}42.^{s}42 −27^{o}50′25′′.4 24.2 0′′.01 W, 0′′.20 S
2002kc Dec. 21.50 3 32 34.72 −27 39 58.3 21.7 0′′.28 W, 0′′.85 S
2002kd Dec. 21.64 3 32 22.34 −27 44 26.9 23.8 3′′.13 W, 0′′.98 S
2002ke Dec. 21.90 3 31 58.77 −27 45 00.7 25.7 1′′.15 E, 0′′.26 S
2002 December 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green